OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 4 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now.

OpenStudy (kropot72):

Let r be the present age of Ralph and s be the present age of Sara. r = 3s ............(1) r + 4 = 2(s + 4) .........(2) Plugging the value of r from equation (1) into equation (2) gives 3s + 4 = 2s + 8 ........(3) Now you need to solve equation (3) to find the value of s. Then plug the value of s into equation (1) to find the value of r.